OpenStudy (anonymous):

Give an example and explain how to apply the quadratic formula to solve a problem and the steps used.

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

take an example : \(x^2+10x+9=0\)